a dog drags its bes box across a floor by applying a horizontal force of 4.5N. the kinetic frictional force acting on the box has magnitude 2.7N. As the box is dragged through 0.76m along the way, what are (a) the work done by the dogs applied force and (b) the increase in thermal energy of the bed and floor?
Answer #1

f=2-7 N > 4.5 N a) Net force Fnet 4.5-2. =1-8N Work done by applied force displacement Fnet = 1-880-76 J = 1 368 J b) Thermal